04. Prove that Var 1/-e + + m te, I in cylindrical coordinates. in upberical coordinates. Orrekin
Answer #1

These relations can be easily proven by using general formula for gradient in orthogonal curvilinear coordinates which is given by:

1 afat Jest hz Jur vf= hou 1 01. én + haus 1 af

Where U1, U2, U: are orthogonal curvilinear coordinates , ei, 62, 63 are the unit vectors in the direction of U1, U2, U:,

and h_1 is the magnitude of tangent vectors to the curve u_1 at a point where u_2 and u_3 are constant i.e hi = Out , similarly др h2 = ди and h3 = диз , here \vec{p} is the position vector of the point with curvilinear

coordinates (u_1,u_2,u_3).

Proof 1:

Let \vec{p} be the position vector of point P, therefore we have \vec{p}=x\hat{i}+y\hat{j}+z\hat{k}

Now as in cylindrical coordinates x=rcos\theta,y=rsin\theta,z=z therefore we have

p=rcosti + rsinaj + zk

Now to apply the above formula we need to find h_1,h_2,h_3 . Which are as follows:

h_1=\left |\frac{\partial \vec{p}}{\partial r}\right |=\sqrt{cos^2\theta+sin^2\theta}=1

h_2=\left |\frac{\partial \vec{p}}{\partial \theta}\right |=\sqrt{r^2(sin^2\theta+cos^2\theta)}=r

h_3=\left |\frac{\partial \vec{p}}{\partial z}\right |=1

and in cylindrical coordinates the unit vectors will be ér, ég, és

Substituting these values in the general formula for gradient in the curvilinear coordinates we get the required result:

Vf - Ofe 10f Ofer + of 3460 + oz.

Proof 2:

Let \vec{p} be the position vector of point P, therefore we have \vec{p}=x\hat{i}+y\hat{j}+z\hat{k}

Now as in spherical polar coordinates x=rsin coso, y = rsino sino, z = rcos therefore we have

\vec{p}=rsin\theta cos\phi \hat{i}+rsin\theta sin\phi\hat{j}+rcos\theta\hat{k}

Now to apply the above formula we need to find h_1,h_2,h_3 . Which are as follows:

21 = sin cosoi + sint sinoj + sincos-o + sin-o sin-o + cose

..h1 = V sin20(cos-o + sin26) + cos26 = 1

rcostcosoi +rcost sinoj-rsinek

:. h2 = r2(cos20(cos + sinºn) + cos20) = 1

h3 = -r sino sinoi +rsin cosoi

:. h3 == r2 sin20 (sin6 + cos26) = rsino

in spherical polar coordinates the unit vectors will be ê, ê, eo

Substituting these values in the general formula for gradient in the curvilinear coordinates we get the required result:

er of lof V = ore tra 1 of + rsino a zº

Please give a up vote if you like the answer.

If you dislike the answer then put your queries, feedback, or suggestion in comments section.

If it doesn't get resolved then down vote the answer.